Hasan buys two kinds of cloth materials for school uniforms, shirt material that costs him ₹50 per metre and trouser material that costs him ₹90 per metre. For every 3 meters of the shirt material he buys 2 metres of the trouser material. He sells the materials at 12% and 10% profit respectively. His total sale is ₹36,600. How much trouser material did he buy? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let 2x m of trouser material and 3x m of shirt material be bought by him Selling price of shirt material per meter =₹50 + 50 ×(12/100) =₹56
Selling price of trouser material per meter =₹90 + 90 × (10/100) =₹99
Total amount of sale =₹36,600
According to the question, (2x×99)+(3x×56)=36600
⇒ 198x + 168x = 36600
⇒ 366x = 36600
⇒x = 36600/366
⇒x=100
Total trouser material he bought = 2x = 2 × 100 = 200 m.

Thanks